OpenStudy (anonymous):

Solve −2(2x + 5) − 3 = −3(x − 1). @horsegirl27

OpenStudy (horsegirl27):

So, first distribute -2 and -3. Do you know how to distribute?

OpenStudy (anonymous):

yes

OpenStudy (horsegirl27):

Ok, so distribute on both sides, and tell me what you get :)

OpenStudy (anonymous):

-4X-10-3 = -3X+1 -4X-13=-3X+1 -4X+3X=1+13X -X=14X X=-14

OpenStudy (anonymous):

then i get -4x+5-3=-3x-1

OpenStudy (horsegirl27):

Next, get x on one side, add 3x to both the left and right side

OpenStudy (anonymous):

-1x+5-3=-1

OpenStudy (horsegirl27):

yay! Now, 5-3?

OpenStudy (anonymous):

-1x+2=-1

OpenStudy (horsegirl27):

Goo, now -2 on both sides?

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

but thats not one of my answer choices

OpenStudy (horsegirl27):

hmmm lets go back through our work

OpenStudy (horsegirl27):

What are your answer choices?

OpenStudy (anonymous):

the equation was suppose to be -4x+(-10)-3=-3x-(-3)

OpenStudy (anonymous):

what do i do after

OpenStudy (horsegirl27):

hmm one sec

OpenStudy (anonymous):

am just going to guess and get it wrong

OpenStudy (horsegirl27):

I am so so so so so sorry!! I would never leave in the middle of helping, but my computer froze